# Photon statistics of resonantly driven spectrally diffusive quantum emitters

### v1

@@ Abstract
Solid-state single-photon sources, such as self-assembled quantum dots and color centers in wide gap crystals, are widely seen as major actors in the emerging quantum technologies Aharonovich16 ; Pelucchi22 ; Couteau23 . They can act like individual atoms trapped in a solid matrix, which enables integration into devices. In turn, these quantum emitters are also sensitive to the solid-state environment, which can manifest itself in decoherence and spectral diffusion (SD) of the emitter transition. The investigation of SD processes is therefore crucial for understanding and improving the performance of solid-state single-photon emitters for applications. In particular, fluctuations of the emitter wavelength limit the number of indistinguishable photons emitted by an artificial atom.
@@
Spectral diffusion is associated with a stochastic process where the random variable $\hbar\omega(t)$ describes the spectral position of the center energy. This varying emission line can be characterized using a variety of experimental techniques, depending on the amplitude and time scale of the fluctuations. When the characteristic time of SD is faster than the inverse count rate, spectroscopy falls short, such that more complex techniques have to be envisaged. Amongst the possible strategies, photon correlation Fourier spectroscopy Brokmann06 ; Coolen07 ; Beyler13 and sub-linewidth filtering Sallen10 ; Abbarchi12 can be used, albeit with intrinsic limitations in either temporal or spectral resolution. Delay-dependent Hong-Ou-Mandel interference can also be performed Thoma16 , but is limited to SPEs emitting close-to-indistinguishable photons, and require to integrate coincidences separately for each time delay, which in many cases can turn out experimentally impractical.
@@
Recently, we have experimentally demonstrated that a combination of resonant laser excitation and photon correlations allows to establish the presence of spectral diffusion as well as some of its characteristics Fournier23PRB . Indeed, this technique simultaneously provides very high spectral and time resolutions. The resonant laser drive converts spectral fluctuations into intensity fluctuations, yielding photon bunching that can be measured through the intensity autocorrelation function $g^{(2)}(\tau)$. It offers the advantage to probe more than 10 orders of magnitude of time scales in a single take, and does not need indistinguishable photons, nor is it limited to zero-phonon-line emission. Such experimental work would benefit from a detailed theoretical investigation of the statistics of the emitted light and its connection with the microscopic process causing SD. Photon correlations are indeed well known as a powerful tool to reveal rich signatures of a large variety of fundamental microscopic phenomena involving quantum emitters, such as single photon emission Michler00 , indistinguishability Santori02 , entanglement Akopian06 , superradiance Jahnke16 , blinking Schroeder21 and phase transitions Fink17 .
@@
In the present work, we investigate the relation between the spectral diffusion process of a quantum emitter and the associated photon statistics under resonant excitation. We consider two emblematic stationary Markovian processes that yield Gaussian inhomogeneous broadening: the Ornstein-Uhlenbeck process and the Gaussian random jump process. While the former depicts a continuous spectral diffusion of some solid-state emitters, the latter models emitters undergoing discrete spectral jumps. Section II describes the theoretical framework of these two spectral diffusion mechanisms. Section III is devoted to the analysis of relevant statistical properties of resonantly scattered light when the emitter is subject to either stochastic process. We identify several experimentally measurable signatures that allow to identify and characterize the diffusion process. Finally, in Section IV we discuss the obtained results and we provide some insights into the spectral diffusion process of B centers in hBN Fournier21 ; Gale22 ; Shevitski19 . The framework and methods developed in the present work can be straightforwardly extended to other SD models and other experimental protocols, as we also discuss in section IV and in the Supplemental Material.
@@
In a wide range of physical systems, continuous diffusion can be well described by a Ornstein-Uhlenbeck (OU) process. The OU process is a Markovian and stationary Gaussian process, and as such, is completely characterized by its correlation function $\langle\omega(t+\tau)\omega(t)\rangle=\Sigma^{2}e^{-\tau/\tau_{\mathrm{SD}}}$, where $\Sigma$ is the standard deviation of the associated Gaussian probability distribution function. Originally derived in the frame of nuclear magnetic resonance by Kubo and Anderson Kubo54 ; Anderson54 , it has been shown to also describe both charge noise and spin noise in self-assembled quantum dots Berthelot06 ; Berthelot10 ; Kuhlmann13 as well as spectral diffusion in other condensed matter systems, such as ions or molecules embedded in solid-state matrices Reilly93 ; DeVoe81 . This process emerges when the light source is coupled to a large ensemble of identical and independent two-level fluctuators (Fig. 1a), and is equivalent to the Wiener model of Brownian motion at short times.
Fig. 1b shows an energy trajectory generated numerically using the stochastic differential equation $d\omega=\omega-dt(\omega-\omega_{0})/\tau_{\mathrm{SD}}+\Sigma\sqrt{t/2\tau_{\mathrm{SD}}}W_{t}$, where $\hbar\omega_{0}$ is the center energy of the inhomogeneous distribution, $\tau_{\mathrm{SD}}$ the spectral diffusion correlation time, and $W_{t}$ is the Wiener process (the continuous-time stochastic process that describes standard Brownian motion Durett ). The numerical integration is performed based on the Euler-Maruyama method Kloeden92 . An example of sample path is shown Fig 1b. The trajectories exhibit a continuous drift such that, after a time larger than the correlation time $\tau_{\mathrm{SD}}$, the energy position is randomly found within a Gaussian envelope of standard deviation $\Sigma$.
@@
Spectral diffusion can also take the form of discrete jumps, occuring at random times governed by a Poisson process, and sampling a Gaussian distribution of standard deviation $\Sigma$. Such stochastic process is termed Gaussian random jump (GRJ) model by Spokoyny et al. Spokoyny20 – we also adopt this terminology in the present work. This process is also a Markovian and stationary process with correlation function $\langle\delta\omega(t)\delta\omega(0)\rangle=\Sigma^{2}e^{-t/\tau_{\mathrm{SD}}}$, and therefore cannot be distinguished from the OU model on the sole basis of its probability density function (PDF) and its correlation function. It has been used to describe spectral fluctuations of nitrogen-vacancy centers in nanodiamonds Wolters13 , molecules in crystalline matrices Ambrose91 , color centers in hBN Spokoyny20 , nitride quantum dots Gao19 , perovskite Utzat19 and CdSe nanocrystals Beyler13 – although, in the latter case, it is the spectral shifts (and not the spectral positions) that sample a Gaussian distribution – which makes little difference at short times, but yields an additional long-time diffusion. Discrete jumps in spectral diffusion typically emerges when an emitter is coupled to a low number of carriers migrating in a manifold of heterogeneous trap states (Fig. 2a), or a low number of neighboring systems whose configuration alternates within a continuum of states, such as ligand configurations in nanocrystals Beyler13 .
@@
In the absence of spectral diffusion, the population of a two-level atom driven by a laser close to resonance reaches a steady-state given by Loudon
@@
In the following, unless explicitly stated, we suppose $\tau_{\mathrm{SD}}\gg T_{1},1/\Sigma$. This corresponds to cases where the spectral fluctuations are slower than both the emitter lifetime and the inverse inhomogeneous linewidth. Most experimental studies of solid-state emitters are consistent with these conditions, at the exception of some recent work demonstrating the achievement of fast SD that yields phenomena such as motional narrowing Berthelot06 ; Berthelot09 ; Pont21 ; Bogaczewicz23 . Out of these specific regimes, a continuously driven emitter is supposed to reach its steady state faster than the spectral fluctuations occur. This justifies the adiabatic approximation, in which the time-dependent intensity writes $C(t)=C_{ss}(\omega(t))$.
@@
where $\langle...\rangle_{t}$ denotes the time averaging, and where quantum effects at $\tau\lesssim T_{1}$ are neglected. As evidenced in prior work Fournier23PRB , resonantly driving a spectrally diffusive emitter leads to bunching, owing to the fact that detection of a photon at time $t_{1}$ informs on a close proximity between the emitter center frequency $\omega(t_{1})$ and the laser frequency $\omega_{L}$. Therefore, the photon detection rate at time $t_{2}\approx t_{1}$ is enhanced with respect to an uncorrelated case where $|t_{2}-t_{1}|\gg\tau_{\mathrm{SD}}$, which translates into bunching.
@@
The short-time bunching can be deduced from the steady-state distribution of the emitter frequencies, independently of the underlying spectral diffusion model Fournier23PRB :
[…]
@@
[…]
@@
where $\omega_{0}$ denotes the center of the inhomogeneous distribution and $\Delta\omega_{\mathrm{inhom}}$ its full width at half maximum ($\Delta\omega_{\mathrm{inhom}}=2\sqrt{2\ln 2}\Sigma$). At zero detuning, we then simply obtain $g^{(2)}(0)=\frac{1}{2\sqrt{\pi\ln 2}}\frac{\Delta\omega_{\mathrm{inhom}}}{\Delta\omega_{\mathrm{hom}}}$. This expression can be interpreted intuitively since, in two-photon correlations, the first detection informs that the spectrally diffusing emitter is close to resonance – which at random times has a likelihood given by the ratio of homogeneous to inhomogeneous linewidth. This leads to an increased likelihood of detecting a second photon, which is maintained until spectral diffusion shifts the emitter spectral position out of resonance. Remarkably, this result shows that it is possible to infer the homogeneous linewidth –and therefore the coherence time of the emitter– from a simple measurement of the inhomogeneous (static) lineshape and the amount of bunching at a fixed laser detuning. Moreover, this estimation does not require any assumptions about the nature of the microscopic mechanism yielding SD. This result is all the more interesting as it is generally difficult to extract information about homogeneous properties in inhomogeneously broadened quantum systems, which can require complex experimental procedures such as spectral hole burning Palinginis03 and dynamical decoupling Abella66 ; Press10 . We note that, while Eq. 5 is derived in the limit of Gaussian inhomogeneous spectra, Eq. 3 provides $\Delta\omega_{\mathrm{hom}}$ in the general case, where any experimental time-averaged lineshape can be accounted for numerically. We further discuss some of these considerations in the Supplemental Material (Sec. S2).
To illustrate the dependence of the short-time bunching on the homogeneous linewidth, on Fig. 5a we plot $B(0)$ as a function of the laser power. The decrease of the amount of bunching as power increases, already observed experimentally Fournier23PRB , is a consequence of power broadening yielding a larger homogeneous response, as discussed in the previous section. The dependence of $B(0)$ on the laser detuning is plotted on Fig. 5a. In this situation, bunching increases with the detuning due to the decreasing likelihood for the emitter to be resonant with the laser when the detuning increases. More examples of calculations are given in the Supplemental Material (Sec. S2). In both cases, we verify numerically that, as expected, the observed behavior is independent of the SD mechanism and correlation time. This is however not the case of the bunching decay, as we expose in the following section.
@@
[…]
$P(\omega^{\prime},t^{\prime}|\omega,t)=\dfrac{\exp\left(-\dfrac{\omega^{\prime}-\omega e^{-(t^{\prime}-t)/\tau_{\mathrm{SD}}}}{2\Sigma^{2}(1-e^{-2(t^{\prime}-t)/\tau_{\mathrm{SD}}})}\right)}{\sqrt{2\pi}\Sigma\sqrt{1-e^{-2(t^{\prime}-t)/\tau_{\mathrm{SD}}}}}$ (7)
@@
Fig. 8 shows the bunching time $\tau_{b}$ as calculated numerically as a function of the laser power. At low powers, it is constant, given by Eq. 10. Above saturation, $\tau_{b}$ increases due to power broadening prolonging the duration of bright periods, and then saturates when the homogeneous linewidth exceeds the width of the inhomogeneous distribution, with in this case $\tau_{b}=\tau_{\mathrm{SD}}/2$. The factor 1/2 originates from the fact that, in the limit $\Delta\omega_{\mathrm{hom}}\gg\Delta\omega_{\mathrm{inhom}}$, the system is first-order insensitive to SD Kuhlmann13 . Therefore, at the second order, fluctuations of $\omega(t)$ at a frequency $f$ yield fluctuations of $C(t)$ at a frequency $2f$ and thus the associated decay time is halved.
@@
[…]
@@
[…]
@@
[…]
@@
We have provided a general framework to describe and analyze the photon statistics of spectrally diffusive emitters resonantly driven by a cw laser. Our simulations, supported by analytical calculations, have evidenced several experimentally accessible signatures of the discrete or continuous nature of the stochastic process governing the variations of the emitter transition energy in time. Two of these characteristic signatures are associated with the intensity correlation function $g^{(2)}$. As a general rule, a decrease in the amount of bunching that is not accompanied by an increase in the bunching decay time is characteristic of discrete jumps to uncorrelated positions.In this regard, a recent experimental study has established the presence of power-dependent bunching in the $g^{(2)}$ of a B center in hBN Fournier23PRB . In light of the results exposed in the present work, we can attribute the SD mechanism of B centers to a discrete jump process, owing to the fact that the bunching timescale is almost independent of the power above saturation (up to $P\approx 100P_{\mathrm{sat}}$). This observation is consistent with what has been demonstrated with other color centers in hBN based on photon correlation Fourier spectroscopy Spokoyny20 . In the latter reference, the GRJ process is evidenced by a direct measurement of the spectral correlation function, which would not be possible in the case of B centers due to their much narrower homogeneous and inhomogeneous widths. As a property of the GRJ model that we have established in section III, the bunching time is equal to the spectral diffusion time, which in the case of B centers has been measured to be $\sim 30$ $\mu$Fournier23PRB . The present work provides guidelines for a more in-depth investigation of spectral diffusion in these systems, as well as in a large variety of other single-photon sources. It can be readily extended to other potential microscopic models, as well as to other experimental processes. As an example, an alternative method that does not require resonant excitation is to filter the emitter fluorescence with a bandwidth narrower than the inhomogeneous linewidth. In the Supplemental Material (Sec. S1), we investigate the statistical properties of light filtered from non-resonantly excited emitters. We show that it shares some similarities with resonance fluorescence, where the spectral selectivity is implemented by the resonant laser. It therefore carries analogous signatures. For instance, observation of a bunching time independent of the filter width implies discrete jumps to the whole inhomogeneous distribution. In addition, the bunching time associated with continuous diffusion is also shorter than the SD time, which possibly modifies the conclusions of prior work using this technique Abbarchi12 . We also present examples of more complex situations, where the emitter subject to blinking (Sec. S4), or where the emitter subject to more than one SD mechanism (Sec. S5) illustrating thereby the versatility of our approach.

### v2

@@ Abstract
Solid-state single-photon sources, such as self-assembled quantum dots and color centers in wide gap crystals, are widely seen as major actors in the emerging quantum technologies [1, 2, 3]. They can act like individual atoms trapped in a solid matrix, which enables integration into devices. In turn, these quantum emitters are also sensitive to the solid-state environment, which can manifest itself in decoherence and spectral diffusion (SD) of the emitter transition. The investigation of SD processes is therefore crucial for understanding and improving the performance of solid-state single-photon emitters for applications. In particular, fluctuations of the emitter wavelength limit the number of indistinguishable photons emitted by an artificial atom.
@@
Spectral diffusion is associated with a stochastic process where the random variable $\hbar\omega(t)$ describes the spectral position of the center energy. This varying emission line can be characterized using a variety of experimental techniques, depending on the amplitude and time scale of the fluctuations. When the characteristic time of SD is faster than the inverse count rate, spectroscopy falls short, such that more complex techniques have to be envisaged. Amongst the possible strategies, photon correlation Fourier spectroscopy [4, 5, 6] and sub-linewidth filtering [7, 8] can be used, albeit with intrinsic limitations in either temporal or spectral resolution. Delay-dependent Hong-Ou-Mandel interference can also be performed [9], but is limited to SPEs emitting close-to-indistinguishable photons, and require to integrate coincidences separately for each time delay, which in many cases can turn out experimentally impractical.
@@
Recently, we have experimentally demonstrated that a combination of resonant laser excitation and photon correlations allows to establish the presence of spectral diffusion as well as some of its characteristics [10]. Indeed, this technique simultaneously provides very high spectral and time resolutions. The resonant laser drive converts spectral fluctuations into intensity fluctuations, yielding photon bunching that can be measured through the intensity autocorrelation function $g^{(2)}(\tau)$. It offers the advantage to probe more than 10 orders of magnitude of time scales in a single take, and does not need indistinguishable photons, nor is it limited to zero-phonon-line emission. Such experimental work would benefit from a detailed theoretical investigation of the statistics of the emitted light and its connection with the microscopic process causing SD. Photon correlations are indeed well known as a powerful tool to reveal rich signatures of a large variety of fundamental microscopic phenomena involving quantum emitters, such as single photon emission [11], indistinguishability [12], entanglement [13], superradiance [14], blinking [15] and phase transitions [16].
@@
In the present work, we investigate the relation between the spectral diffusion process of a quantum emitter and the associated photon statistics under resonant excitation. We consider two emblematic stationary Markovian processes that yield Gaussian inhomogeneous broadening: the Ornstein-Uhlenbeck process and the Gaussian random jump process. While the former depicts a continuous spectral diffusion of some solid-state emitters, the latter models emitters undergoing discrete spectral jumps. Section II describes the theoretical framework of these two spectral diffusion mechanisms. Section III is devoted to the analysis of relevant statistical properties of resonantly scattered light when the emitter is subject to either stochastic process. We identify several experimentally measurable signatures that allow to identify and characterize the diffusion process. Finally, in Section IV we discuss the obtained results and we provide some insights into the spectral diffusion process of B centers in hBN [17, 18, 19]. The framework and methods developed in the present work can be straightforwardly extended to other SD models and other experimental protocols, as we also discuss in section IV and in the Supplemental Material.
@@
In a wide range of physical systems, continuous diffusion can be well described by a Ornstein-Uhlenbeck (OU) process. The OU process is a Markovian and stationary Gaussian process, and as such, is completely characterized by its correlation function $\langle\omega(t+\tau)\omega(t)\rangle=\Sigma^{2}e^{-\tau/\tau_{\mathrm{SD}}}$, where $\Sigma$ is the standard deviation of the associated Gaussian probability distribution function. Originally derived in the frame of nuclear magnetic resonance by Kubo and Anderson [20, 21], it has been shown to also describe both charge noise and spin noise in self-assembled quantum dots [22, 23, 24] as well as spectral diffusion in other condensed matter systems, such as ions or molecules embedded in solid-state matrices [25, 26]. This process emerges when the light source is coupled to a large ensemble of identical and independent two-level fluctuators (Fig. 1a), and is equivalent to the Wiener model of Brownian motion at short times.
Fig. 1b shows an energy trajectory generated numerically using the stochastic differential equation $d\omega=dt(\omega_{0}-\omega)/\tau_{\mathrm{SD}}+\Sigma dW_{t}$, where $\hbar\omega_{0}$ is the center energy of the inhomogeneous distribution, $\tau_{\mathrm{SD}}$ the spectral diffusion correlation time, and $W_{t}$ is the Wiener process (the continuous-time stochastic process that describes standard Brownian motion [27]). The numerical integration is performed based on the Euler-Maruyama method [28]. An example of sample path is shown Fig 1b. The trajectories exhibit a continuous drift such that, after a time larger than the correlation time $\tau_{\mathrm{SD}}$, the energy position is randomly found within a Gaussian envelope of standard deviation $\Sigma$.
@@
Spectral diffusion can also take the form of discrete jumps, occuring at random times governed by a Poisson process, and sampling a Gaussian distribution of standard deviation $\Sigma$. Such stochastic process is termed Gaussian random jump (GRJ) model by Spokoyny et al. [29] – we also adopt this terminology in the present work. This process is also a Markovian and stationary process with correlation function $\langle\delta\omega(t)\delta\omega(0)\rangle=\Sigma^{2}e^{-t/\tau_{\mathrm{SD}}}$, and therefore cannot be distinguished from the OU model on the sole basis of its probability density function (PDF) and its correlation function. It has been used to describe spectral fluctuations of nitrogen-vacancy centers in nanodiamonds [30], molecules in crystalline matrices [31], color centers in hBN [29], nitride quantum dots [32], perovskite [33] and CdSe nanocrystals [6] – although, in the latter case, it is the spectral shifts (and not the spectral positions) that sample a Gaussian distribution – which makes little difference at short times, but yields an additional long-time diffusion. Discrete jumps in spectral diffusion typically emerges when an emitter is coupled to a low number of carriers migrating in a manifold of heterogeneous trap states (Fig. 2a), or a low number of neighboring systems whose configuration alternates within a continuum of states, such as ligand configurations in nanocrystals [6].
@@
In the absence of spectral diffusion, the population of a two-level atom driven by a laser close to resonance reaches a steady-state given by [34]
@@
In the following, unless explicitly stated, we suppose $\tau_{\mathrm{SD}}\gg T_{1},1/\Sigma$. This corresponds to cases where the spectral fluctuations are slower than both the emitter lifetime and the inverse inhomogeneous linewidth. Most experimental studies of solid-state emitters are consistent with these conditions, at the exception of some recent work demonstrating the achievement of fast SD that yields phenomena such as motional narrowing [22, 35, 36, 37]. Out of these specific regimes, a continuously driven emitter is supposed to reach its steady state faster than the spectral fluctuations occur. This justifies the adiabatic approximation, in which the time-dependent intensity writes $C(t)=C_{ss}(\omega(t))$.
@@
where $\langle...\rangle_{t}$ denotes the time averaging, and where quantum effects at $\tau\lesssim T_{1}$ are neglected. As evidenced in prior work [10], resonantly driving a spectrally diffusive emitter leads to bunching, owing to the fact that detection of a photon at time $t_{1}$ informs on a close proximity between the emitter center frequency $\omega(t_{1})$ and the laser frequency $\omega_{L}$. Therefore, the photon detection rate at time $t_{2}\approx t_{1}$ is enhanced with respect to an uncorrelated case where $|t_{2}-t_{1}|\gg\tau_{\mathrm{SD}}$, which translates into bunching.
@@
The short-time bunching can be deduced from the steady-state distribution of the emitter frequencies, independently of the underlying spectral diffusion model [10]:
$g^{(2)}(0)=\dfrac{\langle I^{2}\rangle}{\langle I\rangle^{2}}=\dfrac{\int d\omega\mathcal{P}(\omega)C_{ss}(\omega)^{2}}{\left[\int d\omega\mathcal{P}(\omega)C_{ss}(\omega)\right]^{2}}$ (3)
@@
$g^{(2)}(0)$ $=\dfrac{1}{\mathcal{P}(\omega_{L})}\dfrac{\int d\omega C_{ss}(\omega)^{2}}{\left[\int d\omega C_{ss}(\omega)\right]^{2}}$
@@
where $\omega_{0}$ denotes the center of the inhomogeneous distribution and $\Delta\omega_{\mathrm{inhom}}$ its full width at half maximum ($\Delta\omega_{\mathrm{inhom}}=2\sqrt{2\ln 2}\Sigma$). At zero detuning, we then simply obtain $g^{(2)}(0)=\frac{1}{2\sqrt{\pi\ln 2}}\frac{\Delta\omega_{\mathrm{inhom}}}{\Delta\omega_{\mathrm{hom}}}$. This expression can be interpreted intuitively since, in two-photon correlations, the first detection informs that the spectrally diffusing emitter is close to resonance – which at random times has a likelihood given by the ratio of homogeneous to inhomogeneous linewidth. This leads to an increased likelihood of detecting a second photon, which is maintained until spectral diffusion shifts the emitter spectral position out of resonance. Remarkably, this result shows that it is possible to infer the homogeneous linewidth –and therefore the coherence time of the emitter– from a simple measurement of the inhomogeneous (static) lineshape and the amount of bunching at a fixed laser detuning. Moreover, this estimation does not require any assumptions about the nature of the microscopic mechanism yielding SD. This result is all the more interesting as it is generally difficult to extract information about homogeneous properties in inhomogeneously broadened quantum systems, which can require complex experimental procedures such as spectral hole burning [38] and dynamical decoupling [39, 40]. We note that, while Eq. 5 is derived in the limit of Gaussian inhomogeneous spectra, Eq. 3 provides $\Delta\omega_{\mathrm{hom}}$ in the general case, where any experimental time-averaged lineshape can be accounted for numerically. We further discuss some of these considerations in the Supplemental Material (Sec. S2).
To illustrate the dependence of the short-time bunching on the homogeneous linewidth, on Fig. 5a we plot $B(0)$ as a function of the laser power. The decrease of the amount of bunching as power increases, already observed experimentally [10], is a consequence of power broadening yielding a larger homogeneous response, as discussed in the previous section. The dependence of $B(0)$ on the laser detuning is plotted on Fig. 5a. In this situation, bunching increases with the detuning due to the decreasing likelihood for the emitter to be resonant with the laser when the detuning increases. More examples of calculations are given in the Supplemental Material (Sec. S2). In both cases, we verify numerically that, as expected, the observed behavior is independent of the SD mechanism and correlation time. This is however not the case of the bunching decay, as we expose in the following section.
@@
$g^{(2)}(\tau)=\dfrac{1}{N}\iint d\omega d\omega^{\prime}\mathcal{P}(\omega)C_{ss}(\omega)P(\omega^{\prime},t+\tau|\omega,t)C_{ss}(\omega^{\prime})$ (6)
where $N=[\int\mathcal{P}(\omega)d\omega C_{ss}(\omega)]^{2}$. We have introduced the spectral diffusion kernel [41]
$P(\omega^{\prime},t^{\prime}|\omega,t)=\dfrac{\exp\left(-\dfrac{\left(\omega^{\prime}-\omega e^{-(t^{\prime}-t)/\tau_{\mathrm{SD}}}\right)^{2}}{2\Sigma^{2}(1-e^{-2(t^{\prime}-t)/\tau_{\mathrm{SD}}})}\right)}{\sqrt{2\pi}\Sigma\sqrt{1-e^{-2(t^{\prime}-t)/\tau_{\mathrm{SD}}}}}$ (7)
@@
Fig. 8 shows the bunching time $\tau_{b}$ as calculated numerically as a function of the laser power. At low powers, it is constant, given by Eq. 10. Above saturation, $\tau_{b}$ increases due to power broadening prolonging the duration of bright periods, and then saturates when the homogeneous linewidth exceeds the width of the inhomogeneous distribution, with in this case $\tau_{b}=\tau_{\mathrm{SD}}/2$. The factor 1/2 originates from the fact that, in the limit $\Delta\omega_{\mathrm{hom}}\gg\Delta\omega_{\mathrm{inhom}}$, the system is first-order insensitive to SD [24]. Therefore, at the second order, fluctuations of $\omega(t)$ at a frequency $f$ yield fluctuations of $C(t)$ at a frequency $2f$ and thus the associated decay time is halved.
@@
At a fixed laser detuning, the average intensity –or count rate– that the emitter scatters is simply given by $\langle C(\omega_{L})\rangle=\int d\omega\mathcal{P}(\omega)C(\omega)$. The count rate is typically measured by integrating the photon number over a macroscopic time $T$. This measured intensity fluctuates because of shot noise, with a relative intensity noise given by $\Delta I_{T}/\langle I\rangle_{T}=1/\sqrt{\langle N\rangle_{T}}$, where $\langle N\rangle_{T}$ is the average number of photons detected during T{T}. These Poissonian fluctuations can be mitigated by optimizing the photon collection efficiency. Additional sources of noise can cause the variance $V_{T}$ of the photon number during $T$ to exceed the Poisson limit $V_{T}=\langle N\rangle_{T}$. The Mandel parameter $Q=V_{T}/\langle N\rangle_{T}-1$ relates the long-time intensity fluctuations to the second-order intensity correlation via the following relation [42, 43, 44, 45]
$Q=\dfrac{\langle N\rangle_{T}}{T}\int_{-T}^{T}d\tau\left(1-\dfrac{|\tau|}{T}\right)\left(g^{(2)}(\tau)-1\right)$ (11)
@@
$\dfrac{\Delta I_{T}}{\langle I\rangle_{T}}=\left[\dfrac{1}{T}\int B(\tau)d\tau\right]^{\frac{1}{2}}$ (12)
@@
$\dfrac{\Delta I_{T}}{\langle I\rangle_{T}}=\left[\dfrac{1}{\sqrt{\pi\ln 2}T}\dfrac{\Delta\omega_{\mathrm{inhom}}}{\Delta\omega_{\mathrm{hom}}}\int\tilde{B}(\tau)d\tau\right]^{\frac{1}{2}}$ (13)
@@
We have provided a general framework to describe and analyze the photon statistics of spectrally diffusive emitters resonantly driven by a cw laser. Our simulations, supported by analytical calculations, have evidenced several experimentally accessible signatures of the discrete or continuous nature of the stochastic process governing the variations of the emitter transition energy in time. Two of these characteristic signatures are associated with the intensity correlation function $g^{(2)}$. As a general rule, a decrease in the amount of bunching that is not accompanied by an increase in the bunching decay time is characteristic of discrete jumps to uncorrelated positions.In this regard, a recent experimental study has established the presence of power-dependent bunching in the $g^{(2)}$ of a B center in hBN [10]. In light of the results exposed in the present work, we can attribute the SD mechanism of B centers to a discrete jump process, owing to the fact that the bunching timescale is almost independent of the power above saturation (up to $P\approx 100P_{\mathrm{sat}}$). This observation is consistent with what has been demonstrated with other color centers in hBN based on photon correlation Fourier spectroscopy [29]. In the latter reference, the GRJ process is evidenced by a direct measurement of the spectral correlation function, which would not be possible in the case of B centers due to their much narrower homogeneous and inhomogeneous widths. As a property of the GRJ model that we have established in section III, the bunching time is equal to the spectral diffusion time, which in the case of B centers has been measured to be $\sim 30$ $\mu$[10]. The present work provides guidelines for a more in-depth investigation of spectral diffusion in these systems, as well as in a large variety of other single-photon sources. It can be readily extended to other potential microscopic models, as well as to other experimental processes. As an example, an alternative method that does not require resonant excitation is to filter the emitter fluorescence with a bandwidth narrower than the inhomogeneous linewidth. In the Supplemental Material (Sec. S1), we investigate the statistical properties of light filtered from non-resonantly excited emitters. We show that it shares some similarities with resonance fluorescence, where the spectral selectivity is implemented by the resonant laser. It therefore carries analogous signatures. For instance, observation of a bunching time independent of the filter width implies discrete jumps to the whole inhomogeneous distribution. In addition, the bunching time associated with continuous diffusion is also shorter than the SD time, which possibly modifies the conclusions of prior work using this technique [8]. We also present examples of more complex situations, where the emitter subject to blinking (Sec. S4), or where the emitter subject to more than one SD mechanism (Sec. S5) illustrating thereby the versatility of our approach.